-3x+b=6x for x

-5a+y=-54 for y

Answer by midwood_trail(310)

(1) 3x + b = 6x solve for x
To solve for a variable means to place the letter alone on one side of the equation.
We need to ISOLATE x.
Let's begin by subtracting 3x from both side.
We now have this:
b = 6x - 3x
b = 3x
To find x, divide both sides of the equation by 3.
b/3 = x....This is the value of x and your answer.
Do you see that x is left alone on the right side?
I just solved for x.
=================================================

(2) -5a + y = -54 solve for y
The same applies here.
ISOLATE y: put it all alone on one side of the equation.
To find y, simply add 5a to both sides.
y = 5a - 54
Do you see that y is now alone (or ISOLATED) on the left side of the equation?
That's what it means to solve for a variable, ANY VARIABLE.
Did you follow?